# High-dimensional components of ${\rm SL}(2,\mathbb{C})$-character varieties of prime knots

### v1

@@ 1 Introduction
Within the character variety of a knot, we have special interests in high-dimensional components (HDCs for short), meaning those of dimension >1>1, for several reasons. Firstly, by the Culler-Shalen theory [10], for a 3-manifold $M$, each ideal point of an algebraic curve in $\mathcal{X}(M)$ is associated to an incompressible surface in $M$. If $\dim\mathcal{X}(K)>1$, then a closed incompressible surface in $E_{K}$ is detected (see [7, Proposition 2.4]). Note that if $M$ is a closed 3-manifold obtained by a Dehn filling of $E_{K}$ for such $K$, then $\dim\mathcal{X}(M)>0$. Secondly, the ${\rm SL}(2,\mathbb{C})$ Casson invariant [8] only uses isolated points in the character variety. It will be amazing if HDCs are taken into account. Thirdly, a conjecture suggested by mathematical physics asserts a surprising identity of adjoint Reidemeister torsions [15], and calls for a better understanding about HDC. Fourthly, HDC plays an important role in the study of Kauffman bracket skein module (KBSM) which has become a central object in quantum topology. By [1], if $\dim\mathcal{X}(M)>0$, then the KBSM of $M$, denoted by $\mathcal{S}(M)$, is not finitely generated. Echoing this, recently it was shown [11] that if $M$ is a closed $3$-manifold with $\dim\mathcal{X}(M)>0$, then $\mathcal{S}(M)$ is not “tame” (in the sense of [11, Theorem 3.1 and Proposition 3.2]). In [5] we found a new type of torsion $e$ in $\mathcal{S}(E_{K})$ for a $4$-strand Montesinos knot $K$; the proof of $e\neq 0$ was built on investigating a HDC of $\mathcal{X}(K)$. As a related but more general result, [2, Corollary 2.2] states that if $M$ is closed with $\dim\mathcal{X}(M)>0$, then $\mathcal{S}(M)$ contains torsion.
@@
We give sufficient conditions for a knot $K$ to admit HDCs, meaning that $\mathcal{X}^{\rm irr}(K)$ contains HDCs, and give lower bounds for the dimensions (see Theorem 3.3). Using elementary techniques, we clarify some significant facts that were never mentioned by other people. Then specifically, we improve a result of Paoluzzi and Porti on the dimensions of character varieties of a class of Montesinos knots (see Remark 3.5).
@@
Question 1.1 (Question (7) in [9, Section 9]).
Does there exist a closed atoroidal $3$-manifold $M$ such that $\dim H_{1}(M;\mathbb{Q})\leq 1$ and $\dim\mathcal{X}^{\rm irr}(M)\geq 1$. What if one restricts to 0-surgery on a knot in $S^{3}$?
Based on the experience, we guess that $\dim\mathcal{X}(K)=1$ for any prime knot $K$ with at most $11$ crossings.
@@ 2 Preliminary
Most notations adopted in this paper had been introduced in [3, Section 2] and [4, Section 2]. The reader may refer there if necessary.
We always use bold letters to denote elements of $G$. Let $\mathbf{e}$ denote the identity matrix. For $\mathbf{a},\mathbf{b}\in G$, let $\mathbf{a}\lrcorner\mathbf{b}=\mathbf{a}\mathbf{b}\mathbf{a}^{-1}$.
@@
The following was stated in [5, Lemma 3.1]:
@@
Remark 2.2.
In (ii-2), we can explicitly determine $\mathbf{a}_{3}$ as follows. Let $s_{0}=t^{2}/2-2$. Let $\overline{\mathbf{a}}_{i}=\mathbf{a}_{i}-(t/2)\mathbf{e}$, let $s_{ij}={\rm tr}(\overline{\mathbf{a}}_{i}\overline{\mathbf{a}}_{j})$, and $s_{123}={\rm tr}(\overline{\mathbf{a}}_{1}\overline{\mathbf{a}}_{2}\overline{\mathbf{a}}_{3})$. Note that $s_{ii}=s_{0}$. Then $f_{t}(t_{12},t_{13},t_{23};t_{123})=0$ is equivalent to
$2s_{123}^{2}+\det((s_{ij})_{3\times 3})=0.$
From ${\rm tr}(\overline{\mathbf{a}}_{1}\overline{\mathbf{a}}_{3})=s_{13}$, ${\rm tr}(\overline{\mathbf{a}}_{2}\overline{\mathbf{a}}_{3})=s_{23}$, ${\rm tr}(\overline{\mathbf{a}}_{1}\overline{\mathbf{a}}_{2}\overline{\mathbf{a}}_{3})=s_{123}$ we can solve
$\overline{\mathbf{a}}_{3}=\frac{1}{s_{0}^{2}-s_{12}^{2}}\big((s_{0}s_{13}-s_{12}s_{23})\overline{\mathbf{a}}_{1}+(s_{0}s_{23}-s_{12}s_{13})\overline{\mathbf{a}}_{2}+2s_{123}\overline{\mathbf{a}}_{1}\overline{\mathbf{a}}_{2}-s_{12}s_{123}\mathbf{e}\big).$
@@
To present such $\rho$, it suffices to choose a direction for each arc and label the arc with an element of $G$. For $t\in\mathbb{C}$, let $\mathcal{R}_{t}(K)$ denote the set of $\rho$ such that $\rho(\mathsf{a})\in G(t)$ for each $\mathsf{a}$, and let $\mathcal{R}_{t}^{\rm irr}(K)$ denote the subset of the irreducible ones; let $\mathcal{X}_{t}^{\rm irr}(K)=\mathcal{R}_{t}^{\rm irr}(K)/\sim$.
@@
Given $\rho\in\mathcal{R}(T)$ and $\mathbf{c}\in G$, let $\mathbf{c}\lrcorner\rho\in\mathcal{R}(T)$ send $\mathsf{a}$ to $\mathbf{c}\lrcorner\rho(\mathsf{a})$. Call $\rho,\rho^{\prime}$ conjugate and denote $\rho\sim\rho^{\prime}$ if $\rho^{\prime}=\mathbf{c}\lrcorner\rho$ for some $\mathbf{c}$.
@@
Figure 1: From left to right: a tangle $T$ in $\mathcal{T}^{2}_{2}$; $N(T)$; $D(T)$; $T_{1}\ast T_{2}$; $T_{1}+T_{2}$.
@@
For $p/q\in\mathbb{Q}$, let $[p/q]$ denote the rational tangle associated to $p/q$ (see [3, Section 2.1] for the precise definition). Let $B(p/q)$ denote the rational link $N([p/q])$. Given $p_{1}/q_{1},\ldots,p_{m}/q_{m}\in\mathbb{Q}$, denote $D([p_{1}/q_{1}]\ast\cdots\ast[p_{m}/q_{m}])$ by $M(p_{1}/q_{1},\ldots,p_{m}/q_{m})$ and call it a Montesinos link.
@@
3 High-dimensional components
@@ 3.1 A general structure
Lemma 3.1.
Given $\mathbf{a}\in G(t)$, $r\in\mathbb{B}_{t}$, let $\mathcal{C}_{t}^{r}(\mathbf{a})=\{\mathbf{x}\in G(t)\colon{\rm tr}(\mathbf{a}\mathbf{x})=r\}$. Then $\mathcal{C}_{t}^{r}(\mathbf{a})\cong\mathbb{C}^{\times}$ if $t\neq\pm 2$, and $\mathcal{C}_{t}^{r}(\mathbf{a})\cong\mathbb{C}$ if $t\in\{\pm 2\}$.
Proof.
Let
$\mathbf{d}(\kappa)=\left(\begin{array}[]{cc}\kappa&0\\ 0&\kappa^{-1}\end{array}\right),\qquad\mathbf{p}=\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right).$
Suppose $\mathbf{b}\in\mathcal{C}_{t}^{r}(\mathbf{a})$, with
$\mathbf{b}=\left(\begin{array}[]{cc}t/2+a&b\\ c&t/2-a\end{array}\right),\qquad bc=\frac{t^{2}}{4}-1-a^{2}.$
When $t=2\varepsilon$ with $\varepsilon\in\{\pm 1\}$, up to conjugacy we may assume $\mathbf{a}=\varepsilon\mathbf{p}$. From ${\rm tr}(\mathbf{a}\mathbf{x})=r$ we see $c=\varepsilon(r-2)\neq 0$, and $b$ is determined by $a$ via
$b=-\frac{a^{2}}{c}=\frac{a^{2}}{\varepsilon(2-r)}.$
@@
    If $(\mathbf{x}_{1},\mathbf{b})$ is irreducible, then by Lemma 2.1 and Remark 2.2, there are two or one matrices $\mathbf{x}_{2}\in G(t)$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$ and ${\rm tr}(\mathbf{x}_{2}\mathbf{b})=a_{3}$.
@@
    If $(\mathbf{x}_{1},\mathbf{b})$ is reducible and $\mathbf{x}_{1}\mathbf{b}\neq\mathbf{b}\mathbf{x}_{1}$, then $t\neq\pm 2$, and up to conjugacy we may assume
    $\mathbf{x}_{1}=\left(\begin{array}[]{cc}\kappa&0\\ 0&\kappa^{-1}\end{array}\right),\qquad\mathbf{b}=\left(\begin{array}[]{cc}\kappa^{\varepsilon}&1\\ 0&\kappa^{-\varepsilon}\end{array}\right),$
    with $\kappa+\kappa^{-1}=t$, $\varepsilon\in\{\pm 1\}$. For $\mathbf{x}_{2}=(y_{ij})_{2\times 2}\in G(t)$, the conditions ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$ and ${\rm tr}(\mathbf{b}\mathbf{x}_{2})=a_{3}$ are equivalent to
@@
    If $\mathbf{x}_{1}\mathbf{b}=\mathbf{b}\mathbf{x}_{1}$, so that $\mathbf{x}_{1}=\mathbf{b}^{\varepsilon}$ with $\varepsilon\in\{\pm 1\}$, then the space of $\mathbf{x}_{2}$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$ and ${\rm tr}(\mathbf{b}\mathbf{x}_{2})=a_{3}$ is
    $\begin{cases}\mathcal{C}_{t}^{a_{3}}(\mathbf{b}),&a_{2}=\varepsilon a_{3}+(1-\varepsilon)t^{2}/2\\ \emptyset,&\text{otherwise}\end{cases}.$
By Lemma 3.1, $\dim\mathcal{C}_{t}^{a_{i}}(\mathbf{a})=1$. We have shown $\dim\Phi^{-1}(\mathcal{C}_{t}^{a_{1}}(\mathbf{a})\setminus\{\mathbf{b}^{\pm 1}\})=1$ and $\dim\Phi^{-1}(\mathbf{b}^{\pm 1})\leq 1$. Thus, $\dim\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})=1$.
Suppose $n=k+1\geq 4$ and suppose the assertion holds for $n=k$. For $\underline{a}=(a_{1},\ldots,a_{k+1})$, let $\underline{a}^{\prime}=(a_{1},\ldots,a_{k})$, and define
@@
The pre-image of each $\mathbf{x}_{k}\in\mathcal{C}_{t}^{a_{k+1}}(\mathbf{b})$ is $\mathcal{M}_{t}^{\underline{a}^{\prime}}(\mathbf{a},\mathbf{x}_{k})$, which has dimension k2k-2 by the inductive hypothesis. Thus, $\dim\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})=k-1=n-2$. ∎
@@
Let $\mathbf{a}=\varrho_{0}(T_{0}^{\rm se})$, $\mathbf{b}=\varrho_{0}(T_{0}^{\rm ne})^{-1}$. Let $a_{i}=t^{2}-{\rm tr}_{v}(\varrho_{i})$, $1\leq i\leq n$.
@@
For each $1\leq i\leq n$, since ${\rm tr}_{v}(\varrho_{i})={\rm tr}(\mathbf{x}_{i-1}^{-1}\mathbf{x}_{i})=t^{2}-a_{i}\in\mathbb{B}_{t}$, by Lemma 2.1 (ii-1) there exists $\mathbf{c}_{i}\in G$ such that
@@
Let $\rho_{0}=\varrho_{0}$. Due to the compatibilities of the values at common arcs,
@@
This defines a map $\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})\to\mathcal{R}_{t}^{\rm irr}(K)$.
Suppose $\mathbf{a}\mathbf{b}\neq\mathbf{b}\mathbf{a}$. If $\rho^{\underline{\mathbf{x}}}\sim\rho^{\underline{\mathbf{x}}^{\prime}}$, say $\rho^{\underline{\mathbf{x}}^{\prime}}=\mathbf{c}\lrcorner\rho^{\underline{\mathbf{x}}}$ for some $\mathbf{c}\in G$, then $\mathbf{c}\lrcorner(\mathbf{a},\mathbf{b})=(\mathbf{a},\mathbf{b})$, so $\mathbf{c}=\pm\mathbf{e}$, implying $\underline{\mathbf{x}}=\underline{\mathbf{x}}^{\prime}$. Hence $\dim\mathcal{X}_{t}^{\rm irr}(K)\geq n-2$ for each generic $t$. Thus, $\dim\mathcal{X}^{\rm irr}(K)\geq n-1$.
Now suppose $\mathbf{a}\mathbf{b}=\mathbf{b}\mathbf{a}$, so that $\mathbf{b}\in\{\mathbf{a}^{\pm 1}\}$. Fix $\mathbf{v}\in\mathcal{C}_{t}^{a_{1}}(\mathbf{a})$. For any $\underline{\mathbf{x}}^{\prime}=(\mathbf{x}^{\prime}_{1},\ldots,\mathbf{x}^{\prime}_{n-1})\in\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})$, since ${\rm tr}(\mathbf{a}\mathbf{x}^{\prime}_{1})={\rm tr}(\mathbf{a}\mathbf{v})\in\mathbb{B}_{t}$, by Lemma 2.1 (ii-1) we can find $\mathbf{c}\in G$ such that $\mathbf{c}\lrcorner(\mathbf{a},\mathbf{x}^{\prime}_{1})=(\mathbf{a},\mathbf{v})$. Note that $\mathbf{c}\lrcorner\mathbf{b}=\mathbf{b}$. Hence $\rho^{\underline{\mathbf{x}}^{\prime}}\sim\rho^{\underline{\mathbf{x}}}$ for some $\underline{\mathbf{x}}$ with $\mathbf{x}_{1}=\mathbf{v}$. This shows $\dim\mathcal{X}_{t}^{\rm irr}(K)\geq n-3$ for each generic $t$. Thus, $\dim\mathcal{X}^{\rm irr}(K)\geq n-2$. ∎
@@
According to [13, Section 7] (also see [5, Section 3.1]), for each generic $t$, there exists $\sigma_{i}\in\mathcal{R}^{\rm irr}(B(\beta_{i}/\alpha_{i}))$, which identifies with some $\varrho_{i}\in\mathcal{R}_{t}(T_{i})$ such that $\mathbf{g}_{\varrho_{i}}=\mathbf{e}$ and ${\rm tr}_{v}(\varrho_{i})\neq 2,t^{2}-2$. In particular, $\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})\neq\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})$. The assertion follows from Theorem 3.3. ∎
@@
On restriction to the case $2\mid\alpha_{0}$ (and $2\nmid\alpha_{i}$ for $1\leq i\leq n$), this improves the result of [14, Theorem 1] that $\dim\mathcal{X}(K)\geq n-2$.
For a connected sum of knots, essentially we achieve the same effect as the “bending trick” does. The key difference is that [14, Section 4] requires $\rho(\mu_{1})\rho(\mu_{n+1})=\rho(\mu_{n+1})\rho(\mu_{1})$. The commuting condition decreases the dimension by $1$, as seen from the last paragraph of the proof of $3.$

### v2

@@ 1 Introduction
Within the character variety of a knot, we have special interests in high-dimensional components (HDCs for short), meaning those of dimension >1>1, for several reasons. Firstly, by the Culler-Shalen theory [9], for a 3-manifold $M$, each ideal point of an algebraic curve in $\mathcal{X}(M)$ is associated to an incompressible surface in $M$. If $\dim\mathcal{X}(K)>1$, then a closed incompressible surface in $E_{K}$ is detected (see [6, Proposition 2.4]). Note that if $M$ is a closed 3-manifold obtained by a Dehn filling of $E_{K}$ for such $K$, then $\dim\mathcal{X}(M)>0$. Secondly, the ${\rm SL}(2,\mathbb{C})$ Casson invariant [7] only uses isolated points in the character variety. It will be amazing if HDCs are taken into account. Thirdly, a conjecture suggested by mathematical physics asserts a surprising identity of adjoint Reidemeister torsions [15], and calls for a better understanding of HDC. Fourthly, HDC plays an important role in the study of Kauffman bracket skein module (KBSM) which has become a central object in quantum topology. By [1], if $\dim\mathcal{X}(M)>0$, then the KBSM of $M$, denoted by $\mathcal{S}(M)$, is not finitely generated. Echoing this, recently it was shown [10] that if $M$ is a closed $3$-manifold with $\dim\mathcal{X}(M)>0$, then $\mathcal{S}(M)$ is not “tame” (in the sense of [10, Theorem 3.1 and Proposition 3.2]). In [4] we found a new type of torsion $e$ in $\mathcal{S}(E_{K})$ for a $4$-strand Montesinos knot $K$; the proof of $e\neq 0$ was built on investigating a HDC of $\mathcal{X}(K)$. As a related but more general result, [2, Corollary 2.2] states that if $M$ is closed with $\dim\mathcal{X}(M)>0$, then $\mathcal{S}(M)$ contains torsion.
@@
We give sufficient conditions for a knot $K$ to admit HDCs, meaning that $\mathcal{X}^{\rm irr}(K)$ contains HDCs, and give lower bounds for the dimensions (see Theorem 3.2). Then specifically, we improve a result of Paoluzzi and Porti on the dimensions of character varieties of a class of Montesinos knots (see Remark 3.4). We also find two kinds of “conditional” substructures for $K$ to admit HDCs, in Section 3.2. Along the way, we clarify some useful facts that were never mentioned by other people, using elementary techniques.
@@
Question 1.1 (Question (7) in [8, Section 9]).
Does there exist a closed atoroidal $3$-manifold $M$ such that $\dim H_{1}(M;\mathbb{Q})\leq 1$ and $\dim\mathcal{X}^{\rm irr}(M)\geq 1$? What if one restricts to 0-surgery on a knot in $S^{3}$?
Acknowledgements
I’d like to thank Professor Jiming Ma for inspiring conversations when I visited Fudan University, and thank Doctor Seokbeom Yoon for valuable discussions. I also thank my colleague Professor Yubo Luo for improving the working environment.
@@ 2 Preliminary
Most notations adopted in this paper had been introduced in [3, Section 2]. One may refer there if necessary.
Use bold letters to denote elements of $G$. For $\mathbf{a},\mathbf{b}\in G$, let $\mathbf{a}\lrcorner\mathbf{b}=\mathbf{a}\mathbf{b}\mathbf{a}^{-1}$.
For $\kappa\in\mathbb{C}^{\ast}=\mathbb{C}\setminus\{0\}$, $u\in\mathbb{C}$, put
$\mathbf{d}(\kappa)=\left(\begin{array}[]{cc}\kappa&0\\ 0&\kappa^{-1}\end{array}\right),\qquad\mathbf{u}(\kappa)=\left(\begin{array}[]{cc}\kappa&1\\ 0&\kappa^{-1}\end{array}\right),\qquad\mathbf{p}(u)=\left(\begin{array}[]{cc}1&u\\ 0&1\end{array}\right).$
Let $\mathbf{e}=\mathbf{d}(1)$, the identity matrix, and let $\mathbf{p}=\mathbf{p}(1)$.
@@
Each element of $G(t)$ with $t\neq\pm 2$ is conjugate to $\mathbf{d}(\kappa)$ for some $\kappa$ with $\kappa+\kappa^{-1}=t$, and each element of $G(2\varepsilon)$ with $\varepsilon\in\{\pm 1\}$ is conjugate to $\varepsilon\mathbf{p}$.
For $\mathbf{a}\in G$, let ${\rm Cen}(\mathbf{a})=\{\mathbf{c}\in G\colon\mathbf{c}\mathbf{a}=\mathbf{a}\mathbf{c}\}$, the centralizer of $\mathbf{a}$.
Given $\mathbf{a}\in G(t)$, $r\in\mathbb{B}_{t}$, let $\mathcal{C}_{t}^{r}(\mathbf{a})=\{\mathbf{x}\in G(t)\colon{\rm tr}(\mathbf{a}\mathbf{x})=r\}$.
Lemma 2.1.
(i) For each $\mathbf{a}\in G(t)$, we have ${\rm Cen}(\mathbf{a})\cong\mathbb{C}^{\ast}$ if $t\neq\pm 2$, and ${\rm Cen}(\mathbf{a})\cong\mathbb{C}\sqcup\mathbb{C}$ if $t\in\{\pm 2\}$. In any case,
${\rm Cen}(\mathbf{a})=\{\mu\mathbf{a}+\nu\mathbf{e}\colon\mu^{2}+t\mu\nu+\nu^{2}=1\}.$
(ii) If $\mathbf{a}\mathbf{b}\neq\mathbf{b}\mathbf{a}$, then ${\rm Cen}(\mathbf{a})\cap{\rm Cen}(\mathbf{b})=\{\pm\mathbf{e}\}$.
Proof.
(i) Up to conjugacy we may assume $\mathbf{a}=\mathbf{d}(\kappa)$ with $\kappa\neq\pm 1$ or $\mathbf{a}=\varepsilon\mathbf{p}$ with $\varepsilon\in\{\pm 1\}$. In the former case, ${\rm Cen}(\mathbf{a})=\{\mathbf{d}(\mu)\colon\mu\in\mathbb{C}^{\ast}\}$; in the latter case, ${\rm Cen}(\mathbf{a})=\{\pm\mathbf{p}(u)\colon u\in\mathbb{C}\}$. Then the assertions are easy to verify.
(ii) By (i), each $\mathbf{c}\in{\rm Cen}(\mathbf{a})$ has the form $\mu\mathbf{a}+\nu\mathbf{e}$. If furthermore $\mathbf{c}\in{\rm Cen}(\mathbf{b})$, then $\mu\mathbf{a}\mathbf{b}=\mu\mathbf{b}\mathbf{a}$, hence $\mu=0$, implying $\mathbf{c}\in\{\pm\mathbf{e}\}$. ∎
Lemma 2.2.
For any $\mathbf{a}\in G(t)$, $r\in\mathbb{B}_{t}$, we have $\mathcal{C}_{t}^{r}(\mathbf{a})\cong\mathbb{C}^{\ast}$ if $t\neq\pm 2$, and $\mathcal{C}_{t}^{r}(\mathbf{a})\cong\mathbb{C}$ if $t\in\{\pm 2\}$.
Proof.
Each $\mathbf{x}\in\mathcal{C}_{t}^{r}(\mathbf{a})$ can be written as
$\mathbf{x}=\left(\begin{array}[]{cc}t/2+a&b\\ c&t/2-a\end{array}\right),\qquad bc=\frac{t^{2}}{4}-1-a^{2}.$
When $t=2\varepsilon$ with $\varepsilon\in\{\pm 1\}$, up to conjugacy we may assume $\mathbf{a}=\varepsilon\mathbf{p}$. From ${\rm tr}(\mathbf{a}\mathbf{x})=r$ we see $c=\varepsilon(r-2)\neq 0$, then $b$ is determined by $a$ via $b=-a^{2}/c$. ∎
@@
The following was stated in [4, Lemma 3.1]; there, actually $t=\pm 2$ can be included. One may refer to [11, Section 2 and 5].
@@
To present such $\rho$, it suffices to choose a direction for each arc and label this arc with an element of $G$. For $t\in\mathbb{C}$, let $\mathcal{R}_{t}(K)$ denote the set of $\rho$’s such that $\rho(\mathsf{a})\in G(t)$ for each directed arc $\mathsf{a}$, and let $\mathcal{R}_{t}^{\rm irr}(K)$ denote the subset of the irreducible ones; let $\mathcal{X}_{t}^{\rm irr}(K)=\{\chi_{\rho}\colon\rho\in\mathcal{R}_{t}^{\rm irr}(K)\}$, which identifies with the set of conjugacy classes of $\rho\in\mathcal{R}_{t}^{\rm irr}(K)$.
@@
Given $\rho\in\mathcal{R}(T)$ and $\mathbf{c}\in G$, let $\mathbf{c}\lrcorner\rho\in\mathcal{R}(T)$ send each directed arc $\mathsf{a}$ to $\mathbf{c}\lrcorner\rho(\mathsf{a})$. Call $\rho,\rho^{\prime}$ conjugate and denote $\rho\sim\rho^{\prime}$ if $\rho^{\prime}=\mathbf{c}\lrcorner\rho$ for some $\mathbf{c}$.
@@
Figure 1: From left to right: $T\in\mathcal{T}^{2}_{2}$; $N(T)$; $D(T)$; $T_{1}\ast T_{2}$; $T_{1}+T_{2}$.
@@
A positive (reps. negative) crossing is regarded as a tangle, denoted by $[1]$ (resp. [1][-1]). For $k\neq 0$, the horizontal composite of $|k|$ copies of $[1]$ (resp. [1][-1]) is denoted by $[k]$ if $k>0$ (resp. $k<0$), and the vertical composite of $|k|$ copies of $[1]$ (resp. [1][-1]) is denoted by $[1/k]$ if $k>0$ (resp. $k<0$).
To each $p/q\in\mathbb{Q}$ is associated a rational tangle
$p/q]:=\begin{cases}[k_{1}]\ast[1/k_{2}]+\cdots+[k_{s}],&2\nmid s\\ [1/k_{1}]+[k_{2}]\ast\cdots+[k_{s}],&2\mid s\end{cases},$
if $k_{s};k_{s-1},\ldots,k_{1}]$ is a continued fraction for $p/q$. Let $B(p/q=N([p/q]))$, a rational link. Given $p_{1}/q_{1},\ldots,p_{m}/q_{m}\in\mathbb{Q}$, denote $D([p_{1}/q_{1}]\ast\cdots\ast[p_{m}/q_{m}])$ by $M(p_{1}/q_{1},\ldots,p_{m}/q_{m})$ and call it a Montesinos link.
@@
3 Structures for high-dimensional components
@@ 3.1 A general structure
    If $(\mathbf{x}_{1},\mathbf{b})$ is irreducible, then by Lemma 2.3 (ii-2), there are two or one matrix $\mathbf{x}_{2}\in G(t)$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$ and ${\rm tr}(\mathbf{x}_{2}\mathbf{b})=a_{3}$.
@@
    If $(\mathbf{x}_{1},\mathbf{b})$ is reducible and $\mathbf{x}_{1}\mathbf{b}\neq\mathbf{b}\mathbf{x}_{1}$, then $t\neq\pm 2$, and up to conjugacy we may assume $\mathbf{x}_{1}=\mathbf{d}(\kappa)$, $\mathbf{b}=\mathbf{u}(\kappa^{\varepsilon})$, with $\kappa+\kappa^{-1}=t$ and $\varepsilon\in\{\pm 1\}$. For $\mathbf{x}_{2}=(y_{ij})_{2\times 2}\in G(t)$, the conditions ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$ and ${\rm tr}(\mathbf{b}\mathbf{x}_{2})=a_{3}$ are equivalent to
@@
    If $\mathbf{x}_{1}\mathbf{b}=\mathbf{b}\mathbf{x}_{1}$, so that $\mathbf{x}_{1}=\mathbf{b}^{\varepsilon}$ with $\varepsilon\in\{\pm 1\}$, then it is easy to see that the space of $\mathbf{x}_{2}$ with ${\rm tr}(\mathbf{x}_{1}\mathbf{x}_{2})=a_{2}$ and ${\rm tr}(\mathbf{b}\mathbf{x}_{2})=a_{3}$ is $\mathcal{C}_{t}^{a_{3}}(\mathbf{b})$ if $a_{2}=\varepsilon a_{3}+(1-\varepsilon)t^{2}/2$, and $\emptyset$ otherwise.
By Lemma 2.2, $\dim\mathcal{C}_{t}^{a_{1}}(\mathbf{a})=\dim\mathcal{C}_{t}^{a_{3}}(\mathbf{b})=1$. Therefore, we have shown that $\dim\Phi^{-1}(\mathcal{C}_{t}^{a_{1}}(\mathbf{a})\setminus\{\mathbf{b}^{\pm 1}\})=1$ and $\dim\Phi^{-1}(\mathbf{b}^{\pm 1})\leq 1$, so $\dim\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})=1$.
Now suppose $n=k+1\geq 4$ and that the assertion holds for $n=k$. For $\underline{a}=(a_{1},\ldots,a_{k+1})$, let $\underline{a}^{\prime}=(a_{1},\ldots,a_{k})$, and define
@@
The pre-image of each $\mathbf{x}\in\mathcal{C}_{t}^{a_{k+1}}(\mathbf{b})$ is $\mathcal{M}_{t}^{\underline{a}^{\prime}}(\mathbf{a},\mathbf{x})$, which is (k2)(k-2)-dimensional by the inductive hypothesis. Thus, $\dim\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})=k-1=n-2$. ∎
@@
Let $t\in\mathbb{C}$ be generic; let $\underline{a}=(a_{1},\ldots,a_{n})$, with $a_{i}=t^{2}-{\rm tr}_{v}(\varrho_{i})$. Let $\mathbf{a}=\varrho_{0}(T_{0}^{\rm se})$, $\mathbf{b}=\varrho_{0}(T_{0}^{\rm ne})^{-1}$.
@@
For each $1\leq i\leq n$, since ${\rm tr}(\mathbf{x}_{i-1}^{-1}\mathbf{x}_{i})={\rm tr}((t\mathbf{e}-\mathbf{x}_{i-1})\mathbf{x}_{i})=t^{2}-a_{i}={\rm tr}_{v}(\varrho_{i})$ and $t^{2}-a_{i}\in\mathbb{B}_{t}$, by Lemma 2.3 (ii-1) there exists $\mathbf{c}_{i}\in G$ such that
@@
Let $\rho_{0}=\varrho_{0}$. Due to the compatibilities of the values at common arcs:
@@
Let $H={\rm Cen}(\mathbf{a})\cap{\rm Cen}(\mathbf{b})$, and $\overline{H}=H/\{\pm\mathbf{e}\}$. Then $\overline{H}$ freely acts on $\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})$ by conjugation. The map $\underline{\mathbf{x}}\mapsto\chi_{\rho^{\underline{\mathbf{x}}}}$ factors through the action of $\overline{H}$, to define a map
$\Psi_{t}:\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})/\overline{H}\to\mathcal{X}_{t}^{\rm irr}(K).$
If $\rho^{\underline{\mathbf{x}}}\sim\rho^{\underline{\mathbf{x}}^{\prime}}$, say $\rho^{\underline{\mathbf{x}}^{\prime}}=\mathbf{c}\lrcorner\rho^{\underline{\mathbf{x}}}$ for some $\mathbf{c}\in G$, then $\mathbf{c}\in H$ and $\underline{\mathbf{x}}^{\prime}=\mathbf{c}\lrcorner\underline{\mathbf{x}}^{\prime}$. Hence $\Psi_{t}$ is injective.
If $\mathbf{a}\mathbf{b}\neq\mathbf{b}\mathbf{a}$, then by Lemma 2.1 (ii), $H=\{\pm\mathbf{e}\}$. If $\mathbf{a}\mathbf{b}=\mathbf{b}\mathbf{a}$, then $\mathbf{b}\in\{\mathbf{a}^{\pm 1}\}$, so $H={\rm Cen}(\mathbf{a})$; by Lemma 2.1 (i), $\dim H=1$. Consequently,
$\dim\mathcal{X}_{t}^{\rm irr}(K)\geq\dim\mathcal{M}_{t}^{\underline{a}}(\mathbf{a},\mathbf{b})-\dim H=\begin{cases}n-2,&\mathbf{a}\mathbf{b}\neq\mathbf{b}\mathbf{a}\\ n-3,&\mathbf{a}\mathbf{b}=\mathbf{b}\mathbf{a}\end{cases}.$
Then the lower bound for $\dim\mathcal{X}^{\rm irr}(K)$ follows. ∎
@@
According to [13, Section 7] (also see [4, Section 3.1]), for each generic $t$, there exists $\sigma_{i}\in\mathcal{R}^{\rm irr}(B(\beta_{i}/\alpha_{i}))$, which identifies with some $\varrho_{i}\in\mathcal{R}_{t}(T_{i})$ such that $\mathbf{g}_{\varrho_{i}}=\mathbf{e}$ and ${\rm tr}_{v}(\varrho_{i})\neq 2,t^{2}-2$. In particular, $\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})\neq\varrho_{0}(T_{0}^{\rm ne})\varrho_{0}(T_{0}^{\rm se})$. By Theorem 3.2, $\dim\mathcal{X}^{\rm irr}(K)\geq n-1$. ∎
@@
On restriction to the case $2\mid\alpha_{0}$ (and $2\nmid\alpha_{i}$ for $1\leq i\leq n$), This improves the result of [14, Theorem 1] that $\dim\mathcal{X}(K)\geq n-2$. The key ingredient is that [14, Section 4] requires $\rho(\mu_{1})\rho(\mu_{n+1})=\rho(\mu_{n+1})\rho(\mu_{1})$. The commuting condition decreases the dimension by $1$, as seen in the last paragraph of the proof of Theorem 3.2.